A long solenoid is fabricated. By closely winding a wire of diameter 0.5mm over a cylindrical frame so that successive turns nearly touch each other what would be magnetic field at center of solenoid if it carries a current of 2.5A

Grade:12

1 Answers

Yuvraj Sharma
5 years ago
n=1/10^-3=10^3
B=4π*10^-7*10^3*5
=2π*10^-3T
So,it is our answer 2π*10^-3T
Thank you
T is unit telsa 
1T=10^4gauss